Science Report – Nov 11th

Science Report 11 November 2019
SOL: 1
Crew 215 – Expedition Boomerang Crew
Submitted by Science Officer Andrew Wheeler

Science Dome Operations: Activity in the science dome included preparing for the deployment of the micrometeorite sampling grid and labeling the sample bags.

Green Hab: Green Hab operations have continued. See Green Hab report.

RAM Operations: The RAM Dome has been activated and has been accessed for suit batteries and fuses for Suit 10 and Suit 2. See the Operations report.

EVA: The first EVA for science involved a geological reconnaissance to the southwest of the hab where two locations on wash areas were scouted for the continuing sampling for micrometeorites. These sites – MM5 (518462E 4248949N) and MM6 (519990E 4248036N) appear to be seasonal alluvial fan deposits and the sampling grid has been established at MM5. See EVA Report.

EVA Suit Maintenance: The crew engineer has evaluated Suit 10 and Suit 2 and has identified the likely faults. See Operations Report.
Additional activities: Nil.

Copyright © The Mars Society. All rights reserved. | Main Site